# Adaptive Reward Design for Reinforcement Learning in Complex Robotic Tasks

### v1

@@ V-A Experimental Setup
Environments. The following RL domains are used: the taxi domain from OpenAI Gym [20], and three other domains adapted from [6].
@@
Baselines. We compare the proposed approach with the following methods as baselines: Q-learning for reward machines (QRM) with reward shaping [4], counterfactual experiences for reward machines (CRM) with reward shaping and hierarchical RL for reward machines (HRM) with reward shaping [6]. We use the code accompanying publications.
@@
RL Algorithms. We use DQN [21] for learning in discrete domains (office world and taxi world), DDQN [22] for water world with continuous state space, and DDPG [23] for HalfCheetah with continuous action space. Note that QRM implementation does not work with DDPG, so we only use HRM and CRM as the baselines for HalfCheetah. We also apply PPO [24] and A2C [25] to HalfCheetah (QRM, CRM and HRM baselines are not compatible with these RL algorithms) and report results in Appendix A. Our implementation was built upon OpenAI Stable-Baselines3 [26].
@@ V-B Results Analysis
The significant advantage of our approach is best illustrated in Figure 4, where baselines fail to learn effectively in environments with infeasible tasks. Although baselines apply potential-based reward shaping [7] to assign intermediate rewards, they cannot distinguish between good and bad terminal states (e.g., completing a sub-goal and colliding with an obstacle have the same potential value). In contrast, our approach provides more effective intermediate rewards, encouraging the agent to learn and maximize task progression.
@@
VI Conclusion
We have developed an adaptive reward design approach for complex robotic tasks. This approach uses reward functions designed to incentivize an agent to complete a task specified by a co-safe LTL formula as much as possible and dynamically updates these reward functions during the learning process. Computational experiments demonstrate that our approach is applicable to various discrete and continuous domains and is compatible with a wide range of RL algorithms. Results show that the proposed approach outperforms baselines in terms of faster convergence and higher task completion rates. For the future work, we would like to evaluate the proposed approach on a broader range of robotic tasks and aim to apply it to real-world RL tasks, such as autonomous driving.
@@ Appendix A Correctness
By the definition of adaptive hybrid reward function ${R_{\mathsf{ah},k}^{\otimes}}$ (cf. Equation IV-B), $\eta_{0}\in[0,1]$ and $\eta_{k}=\frac{\eta_{k-1}}{\theta}$ with $\theta>1$. We have $\lim_{k\to\infty}\eta_{k}=0$. The first case of Equation IV-B, $\eta_{k}\cdot-d^{k}_{\varphi}(q)$, tends to 00; and the second case tends to $\max\{\rho^{0}_{\varphi}(q,q^{\prime}),\rho^{k}_{\varphi}(q,q^{\prime})\}$. Thus, it holds that $\lim_{k\to\infty}{R_{\mathsf{ah},k}^{\otimes}}={R_{\mathsf{ap},k}^{\otimes}}$. ∎
@@
Now we are ready to prove Theorem 1 as stated in Section IV and repeated here.
@@
Results for HalfCheetah. Figure 5 shows the results of applying three different RL algorithms, DDPG [23], PPO [24], and A2C [25], to HalfCheetah environments. The comparison between the proposed approach and all baselines using DDPG has already been discussed in Section V. Since the QRM, CRM, and HRM baselines are not compatible with PPO and A2C, we only compare with the naive baseline here.
@@
Comparing different reward functions, we find that the Naive baseline achieves comparable performance with the proposed reward functions in all HalfCheetah environments. However, as noted in Section V, it usually performs the worst in other domains. One possible explanation is that the HalfCheetah task has a unique structure, where each sub-goal requires moving forward by the same distance. The Naive reward function assigns a reward of 1 for each sub-goal, maintaining consistency in the learning process.
@@
Ablation study. Additionally, we conduct an ablation study to investigate the sensitivity of the hyperparameters $\theta$ and $N$ used for updating distance-to-acceptance values (cf. Section IV-B). Figure 6 shows the normalized reward for two infeasible environments: Taxi World and Water World. The results demonstrate that the proposed approach converges with a sufficiently large value of $\theta\in\{2,000,5,000,10,000\}$. Moreover, it takes more training steps to achieve policy convergence with larger values of $N$, indicating longer intervals between consecutive updates of reward functions. Figure 7 shows the success rates for the feasible version of Taxi World and Water World. These results suggest that feasible environments benefit from longer update intervals $N$, as the agent has more time to explore and gather experience before the reward function is modified.

### v2

@@ 5.1 Experimental Setup
Environments. The following RL domains are used: the taxi domain from OpenAI Gym [Brockman et al., 2016], and three other domains adapted from Icarte et al. [2022].
@@
Baselines. We compare the proposed approach with the following methods as baselines: Q-learning for reward machines (QRM) with reward shaping [Camacho et al., 2019], counterfactual experiences for reward machines (CRM) with reward shaping and hierarchical RL for reward machines (HRM) with reward shaping [Icarte et al., 2022]. We also evaluate RM-based algorithms incorporating partial rewards, which are detailed in Appendix C. We use the code accompanying publications.
@@
RL Algorithms. We use DQN Mnih et al. [2015] for learning in discrete domains (office world and taxi world), DDQN [Van Hasselt et al., 2016] for water world with continuous state space, and DDPG [Lillicrap et al., 2016] for HalfCheetah with continuous action space. Note that QRM implementation does not work with DDPG, so we only use HRM and CRM as the baselines for HalfCheetah. We also apply PPO [Schulman et al., 2017] and A2C [Mnih et al., 2016] to HalfCheetah (QRM, CRM and HRM baselines are not compatible with these RL algorithms) and report results in Appendix B due to the page limit. Our implementation was built upon OpenAI Stable-Baselines3 [Raffin et al., 2021].
@@ 5.2 Results Analysis
The significant advantage of our approach is best illustrated in Figure 4, where baselines fail to learn effectively in environments with infeasible tasks. Although baselines apply potential-based reward shaping [Ng et al., 1999] to assign intermediate rewards, they cannot distinguish between good and bad terminal states (e.g., completing a sub-goal and colliding with an obstacle have the same potential value). In contrast, our approach provides more effective intermediate rewards, encouraging the agent to learn and maximize task progression.
@@
Ablation study. Additionally, we conduct an ablation study to investigate the sensitivity of the hyperparameters $\theta$ and $N$ used for updating distance-to-acceptance values (cf. Section 4.2). Figure 5 shows the normalized reward for two infeasible environments: Taxi World and Water World. The results demonstrate that the proposed approach converges with a sufficiently large value of $\theta\in\{2,000,5,000,10,000\}$. Moreover, it takes more training steps to achieve policy convergence with larger values of $N$, indicating longer intervals between consecutive updates of reward functions. Figure 6 shows the success rates for the feasible version of Taxi World and Water World. These results suggest that feasible environments benefit from longer update intervals $N$, as the agent has more time to explore and gather experience before the reward function is modified.
Hyperparameter Selection: Practical Guidance. We offer the following heuristics for selecting key hyperparameters in our framework. For the reward update interval $N$, a useful starting point is the total training budget divided by the number of distinct task stages (e.g., states in a task-governing DFA), as this aims to provide the agent with sufficient interaction episodes within each task stage before potential reward adjustments. The reward scaling factor $\theta$ can be initially set to the sum of progression rewards, $\sum_{q,q^{\prime}}p_{\phi}\left(q,q^{\prime}\right)$, which approximates the cumulative effort. Insights from the ablation study further suggest that task feasibility can guide these choices: potentially infeasible tasks may benefit from smaller $\theta$ values and more frequent updates (smaller $N$) to enable regular progress assessment and dynamic reward adjustment in challenging settings. In contrast, feasible tasks often accommodate larger $\theta$ and $N$ to allow for the collection of more meaningful data before reward function modification. When employing hybrid reward functions, we recommend small magnitudes for step-wise penalties (e.g., $10^{-3}$ or $10^{-4}$) to avoid overwhelming the positive shaping signals. These guidelines serve as practical starting points, though optimal settings can be task-dependent.
6 Conclusion
We have developed a logic-based adaptive reward shaping approach for RL. Our approach uses reward functions designed to incentivize an agent to complete a task specified by a co-safe LTL formula as much as possible, and dynamically updates these reward functions during the learning process. This dynamic reward shaping is beneficial for scenarios where environmental uncertainties can lead to task failure despite successful subtask progress.
Computational experiments demonstrate that our approach is applicable to various discrete and continuous RL domains and is compatible with a wide range of RL algorithms such as DQN, DDQN, DDPG, PPO, and A2C. Experimental results also show that the proposed approach generally outperforms state-of-the-art baselines, achieving faster convergence to a better policy with higher expected return and task completion rate.
There are several directions for future work. First, we will evaluate the proposed approach on a broader range of RL domains beyond the benchmarks used in our experiments. Second, we will explore extending the approach to multi-agent RL. Finally, we aim to apply the proposed approach to real-world RL tasks, such as autonomous driving.
Acknowledgments
This work was supported in part by the U.S. National Science Foundation under Grants CCF-1942836 and CCF-2131511. The opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsoring agencies.
Supplementary Material
@@ Appendix A Correctness
By the definition of adaptive hybrid reward function ${R_{\mathsf{ah},k}^{\otimes}}$ (cf. Equation 4.2), $\eta_{0}\in[0,1]$ and $\eta_{k}=\frac{\eta_{k-1}}{\theta}$ with $\theta>1$. We have $\lim_{k\to\infty}\eta_{k}=0$. The first case of Equation 4.2, $\eta_{k}\cdot-d^{k}_{\varphi}(q)$, tends to 00; and the second case tends to $\max\{\rho^{0}_{\varphi}(q,q^{\prime}),\rho^{k}_{\varphi}(q,q^{\prime})\}$. Thus, it holds that $\lim_{k\to\infty}{R_{\mathsf{ah},k}^{\otimes}}={R_{\mathsf{ap},k}^{\otimes}}$. ∎
@@
Now we are ready to prove Theorem 1 as stated in Section 4 and repeated here.
@@
Appendix B Compatibility with On-Policy Learning
Results for HalfCheetah. Figure 7 shows the results of applying three different RL algorithms, DDPG [Lillicrap et al., 2016], PPO [Schulman et al., 2017], and A2C [Mnih et al., 2016], to HalfCheetah environments. The comparison between the proposed approach and all baselines using DDPG has already been discussed in Section 5. Since the QRM, CRM, and HRM baselines are not compatible with PPO and A2C, we only compare with the naive baseline here.
@@ Appendix B Compatibility with On-Policy Learning
Comparing different reward functions, we find that the Naive baseline achieves comparable performance with the proposed reward functions in all HalfCheetah environments. However, as noted in Section 5, it usually performs the worst in other domains. One possible explanation is that the HalfCheetah task has a unique structure, where each sub-goal requires moving forward by the same distance. The Naive reward function assigns a reward of 1 for each sub-goal, maintaining consistency in the learning process.
@@
Appendix C Partial Rewards in Reward Machines
This ablation study investigates the effect of incorporating partial rewards into Reward Machine (RM) structures along with the use of potential-based reward shaping. While RMs are theoretically capable of representing and utilizing partial rewards (e.g., in the Office World environment, the RM transitions through states $u_{0}$ [initial], $u_{1},u_{2}$ [intermediate], and $u_{3}$ [goal], as depicted in Figure 2 (b) Icarte et al. [2022], our empirical evaluation reveals that their inclusion does not consistently enhance performance and can lead to performance degradation.
To evaluate the impact of partial rewards on RM-based algorithms, we conducted experiments in deterministic environments: Office World and Taxi World. We define "Partial Reward Q-learning Reward Machine" (PR QRM) as the QRM algorithm variant that incorporates partial rewards. For consistency across algorithms, we similarly introduce PR CRM and PR HRM, denoting CRM and HRM variants also utilizing partial rewards. All baselines in this study are equipped with potential-based reward shaping. Across both Office World and Taxi World environments, all RM-based algorithms suffered a performance degradation when supplemented with partial rewards of $1$ for each intermediate step.
These findings suggest that the algorithms, particularly in their current configurations, may not be inherently designed to effectively leverage partial rewards in conjunction with potential reward shaping. One plausible explanation for the observed performance degradation is that, as discussed in Icarte et al. [2022], potential reward shaping can assign positive rewards to actions that lead to undesirable "violation" states within the RM, potentially exacerbating the negative effects of partial rewards. Therefore, careful consideration and potentially algorithm modifications are necessary to effectively harness the benefits of partial rewards within Reward Machine frameworks, especially when integrated with reward shaping techniques. In contrast to these observations, our proposed algorithms are designed to effectively incorporate partial rewards across diverse environments without performance degradation, while ensuring both task completion and reward maximization.
Figure 8: Results in Office World and Taxi World for RM algorithms using partial rewards.